OpenStudy (anonymous):

The circular bull’s-eye of an archery target has a diameter of 24 cm, which is surrounded by 4 concentric rings with a width of 12 cm each. Draw the target in a coordinate plane with center at the origin. Write the equations of the circles that form the boundaries of the different regions of the target.

OpenStudy (anonymous):

I don't understand this at all!

OpenStudy (anonymous):

do you just need the equations of 5 circles? Because that's not too bad, since they're all concentric circles and all

OpenStudy (anonymous):

If that's all it really is, you answer will be the following: x^2+y^2=12^2 x^2+y^2=(24)^2 x^2+y^2=36^2 x^2+y^2=48^2 and finally x^2+y^2=60^2

OpenStudy (anonymous):

Oh ok, I was making it more difficult than I had to. Thanks!